Determine the number of bijective functions f from {1, 2, . . . , 8} to itself such that f(i) 6= i for any even number i.


1
Expert's answer
2021-10-25T03:24:38-0400

the number of bijective functions f from {1, 2, . . . , 8} to itself such that f(i) "\neq" i for any even number i.


The total number of bijective functions:

"N=8!=40320"  

There are 4 even numbers in the set, so the number of bijective functions such that f(i)=i:

 "n=(8-4)!=24"


Answer:

"N-n=40320-24=40296"